# A New Construction Structure on Multi-access Coded Caching with Linear Subpacketization: Cyclic Multi-Access Non-Half-Sum Disjoint Packing

### v2

@@ Abstract
We consider the multi-access coded caching (MACC) system introduced by Hachem et al., which consists of a central server with $N$ files and $K$ cache nodes, each of memory size $M$, where each user can access $L$ cache nodes in a cyclic wrap-around fashion. At present, several existing schemes achieve competitive transmission performance, but their subpacketization levels grow exponentially with the number of users. In contrast, schemes with linear or polynomial subpacketization always incur higher transmission loads. We aim to design a multi-access coded caching scheme with linear subpacketization $F$ while maintaining low transmission load. Recently, Cheng et al. proposed a construction framework for coded caching schemes with linear subpacketization (i.e., $F=K$) called non-half-sum disjoint packing (NHSDP). Inspired by this structure, we introduce a novel combinatorial structure named cyclic multi-access non-half-sum disjoint packing (CMA-NHSDP) by extending NHSDP to MACC system. By constructing CMA-NHSDP, we obtain a new class of multi-access coded caching schemes. Theoretical and numerical analyses show that our scheme achieves lower transmission loads than some existing schemes with linear subpacketization. Moreover, the proposed scheme achievesa lower transmission load compared to existing schemes with exponential subpacketization in some cases.
@@ I Introduction
With the popularity of Internet-connected devices, wireless networks have faced increasing traffic pressure during peak service hours. Furthermore, the highly time-varying nature of network traffic often leads to congestion during peak hours and underutilized resources during off-peak hours. Caching is an effective technique to solute this problem, which pre-stores popular contents at network cache nodes during off-peak periods and exploits local cache resources to obtain “local caching gain”, thereby reducing the overall network traffic[1]. Maddah-Ali and Niesen (MN) [2] proposed the first coded caching scheme, which achieves a “global caching gain” by designing each broadcast message to simultaneously send multiple users.
In the $(K,M,N)$ centralized caching system introduced by Maddah-Ali and Niesen, a central server storing $N$ files of equal size is connected to $K$ users, each equipped with a local cache of size $M$ files. The server communicates with the users over a error-free shared-link. A $(K,M,N)$ coded caching scheme consists of two phases: the placement phase and the delivery phase. In the placement phase, each file is divided into $F$ equal-length data packets, some of which are stored in each user’s cache without knowledge of future demand. The quantity $F$ is referred to as the subpacketization. In the delivery phase, each user randomly requests a file from the server. Based on these requests and the cached content of each user, the server broadcasts coded messages via an error-free shared link to ensure that each user can recover the requested file. The transmission rate $R$ is defined as the worst-case normalized transmission load across all possible user demands, and the design objective of the scheme is to minimize $R$.
The MN scheme employs an uncoded placement strategy [^1] Original files are stored in the user cache directly; otherwise, it would be a coded placement strategy. combined with a coded delivery strategy. Under the constraint of uncoded placement, it has been shown that the MN scheme achieves the optimal rate when $K\leq N$ [3, 4]. For the case $K>N$, Yu et al. [5] further optimized the scheme by eliminating redundant transmissions in the MN scheme, generating an optimal uncoded placement scheme in that regime. Therefore, the MN scheme has been extensively adapted to various network topologies, including but not limited to cache-aided relay networks [6], heterogeneous networks [7], multi-access networks with cyclic wrap-around topology [8].
@@ I-A Multi-access Coded Caching
Under the constraint of uncoded cache placement, the authors of [8] introduced a multi-access coded caching scheme. They proved that the gap between its achievable rate and the information-theoretic lower bound is no more than $cL$, where $c$ is a constant. Reference [22] proposes a new scheme using index coding, which achieves a lower rate than [8] when $K<\frac{KLM}{N}+L$. Furthermore, for the case where $L\geq K/2$ and $N\geq K$, the authors of [22] derived a converse bound under uncoded placement. They proved that the scheme has order optimality within a factor of 2. A transformation method is proposed in [23] to adapt the MN scheme to the multi-access scenario. This adaptation achieves a lower rate than [8] when $L\nmid K$. The work in [24] decomposed the multi-access caching problem into structured index coding problem. Based on solutions to the structured index coding problem, a new scheme has been proposed with with a rate that does not exceed the rates of the schemes in [22] and [23]. It is worth noting that subpacketization of the above scheme increases exponentially with KK.[25] proposed a consecutive cyclic placement strategy. By constructing a class of PDAs under this strategy, a novel multi-access coded caching scheme was introduced. This scheme have a subpacketization increasing linearly with the number of users and attains a minimal rate of $\frac{K-KML/N}{2\left\lfloor\frac{K}{K-KML/N+1}\right\rfloor+1}$. All the aforementioned schemes are designed for general system parameters. A summary is provided in Table I. For specific parameter configurations, several schemes with reduced subpacketization have been introduced in [26, 27, 28, 29]. These are also included in Table I.
TABLE I: Known multi-access coded caching schemes with uncoded placement.Where $K$ is number of users, number of cache nodes is equal to $K$, $L$ is the access degree and the cache-node memory ratio is $\frac{t}{K}$. If the range of $t$ is not specified in the Limitation section, then $t\in\{0,\ldots,\lfloor\frac{K}{L}\rfloor\}$.
@@
$\frac{1}{\left(\lceil\frac{2tL}{K-tL+1}\rceil+1\right)}+\sum_{r=-\frac{K-tL+3}{2}}^{K-tL}\frac{2}{1+\lceil\frac{tL}{r}\rceil}$ $K-tL>1\text{ is odd};\text{$t$, $K$ are coprime}$
@@
Furthermore, the work in [30] introduces a multi-access coded caching scheme that incorporates secure delivery mechanisms. When the additional memory allocated for security purposes is excluded, this scheme achieves a same transmission rate as the RK1 scheme presented in [22], and the required subpacketization level does not exceed $K$. In [31], a converse bound based on the cut-set argument was established, and a new scheme of coded placement was proposed for the memory regime satisfying M<(NK+L)/KM<(N-K+L)/K. This scheme is proven to be optimal under the condition $N\leq K$. Several studies [32, 33, 34] have investigated multi-access coded caching systems with a slightly modified topology based on cross-resolvable designs. The authors of [35, 36] explored a two-dimensional multi-access scenario, where the mobile users move on a two-dimensional grid and can access the nearest cache-nodes inside the grid. The work in [37] presents a novel scheme for a general MACC setting, unifying existing MACC schemes with the customized solution for specific encoding cache scenarios. This scheme enhances user distribution density while maintaining good load performance.
In Table I, it can be observed that while the schemes in [8, 22, 23, 24] achieve competitive transmission rates, their subpacketization grows exponentially with the number of users. In contrast, [25, 28, 29] and other schemes maintain linear or polynomial subpacketization, but at the cost of limited coded caching gains and higher transmission loads. This trade-off reveals a key limitation in multi-access coded caching. Although the PDA serves as a combinatorial structure that characterizes placement and delivery strategies, its definition alone does not directly yield constructions that simultaneously achieve low subpacketization and significant coded caching gain. Therefore, constructing a general framework for PDA that guarantee linear subpacketization while maintaining near-optimal transmission performance is an important and meaningful research direction.
@@ I-B Research motivation and contribution
In this paper, we focus on designing multi-access coded caching schemes with linear subpacketization, aiming to minimum the transmission rate under uncoded placement and one-shot delivery strategies. Through the analysis of Table I, we can find that existing multi-access coded caching schemes with linear or polynomial subpacketization still face challenges in balancing subpacketization and rate. Motivated by the structure of non-half-sum disjoint packing (NHSDP) in shared-link caching systems, we extend this interesting structure to the multi-access network and propose a novel structure called cyclic multi-access non-half-sum disjoint packing (CMA-NHSDP). This structure offers a unified combinatorial framework for designing multi-access coded caching schemes with linear subpacketization. More specifically, the main contributions are summarized as follows:
@@
  • We propose a novel combinational structure termed CMA-NHSDP. It is worth noting that the NHSDP proposed in [19] for shared-link networks cannot be directly applied to MACC systems, as each user in the shared-link system is equipped with independent caches. The structure CMA-NHSDP is the generalized framework for extending NHSDP to MACC systems. Given an $(L,v,g,b)$ CMA-NHSDP for any odd positive integers $v$, any positive integers $L$, $g$ and $b$, we can obtain an $L$-continuous (K=v,F=v,Z=vbg,S=bv)(K=v,F=v,Z=v-bg,S=bv) PDA. When $K$ is even, we can add one virtual user to construct a scheme based on the $(L,K+1,g,b)$ CMA-NHSDP.
@@ A $F$-division $(K,L,M,N)$ multi-access coded caching scheme contains two independent phases:
The transmission rate (or rate) is defined as the worst case transmission amount $R=\max_{{\bf d}\in[1:N]^{K}}S_{\mathbf{d}}/F$. Our objective is to design a scheme such that the transmission load is as small as possible given the parameters $K$, $L$, $M$ and $N$.
@@
  • An $F\times K$ node-placement array $\mathbf{C}$ consists of star and null, where $F$ and $K$ represent the subpacketization and the number of cache-nodes, respectively. The entry located at the position $(j,k)$ in $\mathbf{C}$ is star if and only if the $k^{\text{th}}$ cache-node caches the $j^{\text{th}}$ packet of each $\mathbf{w}_{n}$ where $n\in[N]$.
  • An $F\times K$ user-retrieve array $\mathbf{U}$ consists of star and null, where $F$ and $K$ represent the subpacketization and the number of users, respectively. The entry at the position $(j,k)$ in $\mathbf{U}$ is star if and only if the $k^{\text{th}}$ user can retrieve the $j^{\text{th}}$ packet of each $\mathbf{w}_{n}$ where $n\in[N]$.
@@ II-B Placement delivery array
The authors in [23] also pointed out that if the user-delivery array $\mathbf{Q}$ satisfies the conditions C22 and C33, each user can decode its required file. That is the following statement. Specifically, if $\mathbf{Q}(j,k)=s$ is an integer, it implies that the $j^{\text{th}}$ packet of all files is not accessible to user $U_{k}$, and the server broadcasts a multicast message (i.e., the XOR of all the requested packets indicated by $s$) to all the users at time slot $s$. Condition C33 of Definition 2 ensures that each user can get its desired packet, since all the other packets in the multicast message are accessible to it. The occurrence number of integer $s$ in $\mathbf{Q}$, denoted by $g_{s}$, is called the coded caching gain of the PDA in time slot $s$, since the message broadcast at time slot $s$ can serve $g_{s}$ users simultaneously. Condition C2 of Definition 2 implies that the number of messages broadcasted by the server is exactly $S$, so the rate is $R=S/F$.
@@
    We can check that all the integers in $\mathbf{Q}$ satisfy the Conditions C2 and C3 in Definition 2. According to Definition 1, we can obtain the following delivery strategy by $\mathbf{Q}$.
@@ III Cyclic Multi-access Non-half-sum disjoint packing
We can find that the stars occurs exactly $Z=v-bg=27-3\times 4=15$ times in each column of $\mathbf{P}$, and appears in each row as $L=3$ cyclic consecutive. So the conditions C22 and C44 hold. In addition, there are exactly $S=81$ vectors in $\mathbf{P}$, and each vector occurs at most once in each row and each column. i.e., the conditions C22 and C33-a of Definition 2 hold. Finally, let us consider condition C33-b. Let us first consider the entries $\mathbf{P}(5,0)=\mathbf{P}0,5)=(5,3)$. We have $\mathbf{P}(5,5)=\mathbf{P}(0,0)=*$ which satisfies condition C33-b. Similarly, we can check that all the vectors in $\mathbf{P}$ satisfy condition C33-b) of Definition 2. Therefore, $\mathbf{P}$ is an $L$-continuous $(27,27,15,81)$ PDA that can realize a MACC scheme with memory ratio $\frac{M}{N}=\frac{5}{27}$, subpacketization $F=K=27$, and load $R=3$.
@@ IV Constructions of PDAs via CMA-NHSDP
By the above equation, $\mathfrak{D}$ has $m_{1}m_{2}m_{3}=3$ blocks and the intersection of any blocks is empty. The minimum and maximum values of $\mathfrak{D}$ are 40-40 and $40$, respectively. To ensure that these two sets $[40:0]$ and $[1:40]$ do not have overlap in $\mathbb{Z}_{v}$, the value of $v$ must satisfy $v\geq 2\times 40+1=81$, since 0 is also in $\mathbb{Z}_{v}$. Next, let we verify that $(\mathbb{Z}_{81},\mathfrak{D})$ is a $(3,81,8,3)$ CMA-NHSDP. Now let we consider the half-sum of $38$ and $16$, i.e., $\frac{38+16}{2}=27$, which is not in any block of $\mathfrak{D}$. Similarly, we can check the half-sum of any two different integers in each block of $\mathfrak{D}$. Finally, we can check that the sorted set $\mathscr{D}=\{\pm 14,\pm 15,\pm 16,\pm 20,\pm 21,\pm 22,\pm 32,\pm 33,\pm 34,\pm 38,\pm 39,\pm 40\}$ satisfies that its complement in $\mathbb{Z}_{7}$ can be partitioned into groups of $3$ cyclic consecutive integers, such as $\{17,18,19\}$.
@@
Next, given the parameters $v$ and $n$ (recall that $K=v$ and $g=2^{n}$), we consider how to select $m_{1},\ldots,m_{n}$ such that the memory ratio is minimized. That is, we aim to find the MACC scheme with the minimum required memory ratio while fixing the coded cache gain. Considering the constraint $v\geq 2\phi(m_{1},m_{2},\ldots,m_{n})+1$ into the optimization problem for the selection of $m_{1},m_{2},\ldots,m_{n}$, we have
@@
Since this equation (15) involves modulo arithmetic, it can not be solved directly. However, through abundant computer calculations, we have found that when $m_{i}$ satisfies the following relationship, a good solution can be obtained.
@@
Next, we compare the performance of Theorem 4 and Construction 2 via a concrete example. Specifically, we set $n=4$, $L=3$, and let $m_{2}$ vary in 1,2,3,4{1,2,3,4}. The corresponding parameters derived from the two approaches are summarized in Fig. 2. It can be observed that Theorem 4 achieves a smaller value of $v$ while maintaining the same objective function value $f$.
Figure 2: $v$-$f$ tradeoff comparison between Theorem 4 and Construction 2

### v3

@@ Abstract
We consider the $(K,L,M,N)$ multi-access coded caching system introduced by Hachem et al., which consists of a central server with $N$ files, $K$ cache nodes and $K$ cacheless users, where each user can access $L$ cache nodes in a cyclic wrap-around fashion. At present, several existing schemes achieve competitive transmission load, but their subpacketization levels grow exponentially with the number of users. In contrast, schemes with linear or polynomial subpacketization always incur higher transmission loads. We aim to design a multi-access coded caching scheme with linear subpacketization $F$ while maintaining low transmission load. Recently, Cheng et al. proposed a construction framework for coded caching schemes with linear subpacketization (i.e., $F=K$) called non-half-sum disjoint packing (NHSDP). Inspired by this structure, we introduce a novel combinatorial structure named cyclic multi-access non-half-sum disjoint packing (CMA-NHSDP) by extending NHSDP to MACC system. By constructing CMA-NHSDP, we obtain a new class of multi-access coded caching schemes. Theoretical and numerical analyses show that our scheme achieves lower transmission loads than some existing schemes with linear subpacketization. Moreover, the proposed schemes achieves lower transmission load compared to existing schemes with exponential subpacketization in some case.
@@ I Introduction
With the popularity of Internet-connected devices, wireless networks have faced increasing traffic pressure during peak service hours. Furthermore, the highly time-varying nature of network traffic often leads to congestion during peak hours and underutilized resources during off-peak hours. Caching is an effective technique to solve this problem, which pre-stores popular contents at network cache nodes during off-peak periods and exploits local cache resources to obtain “local caching gain”, thereby reducing the overall network traffic[1]. Maddah-Ali and Niesen (MN) [2] proposed the first coded caching scheme, which achieves a “global caching gain” by designing each broadcast message to simultaneously send multiple users.
In the $(K,M,N)$ centralized caching system introduced by Maddah-Ali and Niesen, a central server storing $N$ files of equal size is connected to $K$ users, each equipped with a local cache of size $M$ files. The server communicates with the users over a error-free shared-link. A $(K,M,N)$ coded caching scheme consists of two phases: the placement phase and the delivery phase. In the placement phase, each file is divided into $F$ equal-length data packets, some of which are stored in each user’s cache without knowledge of future demand. The quantity $F$ is referred to as the subpacketization. In the delivery phase, each user randomly requests a file from the server. Based on these requests and the cached content of each user, the server broadcasts coded messages via an error-free shared link to ensure that each user can recover the requested file. The transmission load $R$ is defined as the worst-case normalized transmission load across all possible user demands, and the design objective of the scheme is to minimize $R$.
The MN scheme employs an uncoded placement strategy [^1] Original files are stored in the user cache directly; otherwise, it would be a coded placement strategy. combined with a coded delivery strategy. Under the constraint of uncoded placement, it has been shown that the MN scheme achieves the optimal transmission load when $K\leq N$ [3, 4]. For the case $K>N$, Yu et al. [5] further optimized the scheme by eliminating redundant transmissions in the MN scheme, generating an optimal uncoded placement scheme in that regime. Therefore, the MN scheme has been extensively adapted to various network topologies, including but not limited to cache-aided relay networks [6], heterogeneous networks [7], multi-access networks with cyclic wrap-around topology [8].
@@ I-A Multi-access Coded Caching
Under the constraint of uncoded cache placement, the authors of [8] introduced a multi-access coded caching scheme. They proved that the gap between its achievable transmission load and the information-theoretic lower bound is no more than $cL$, where $c$ is a constant. Reference [22] proposed a new scheme using index coding, which achieves a lower transmission load than [8] when $K<\frac{KLM}{N}+L$. Furthermore, for the case where $L\geq K/2$ and $N\geq K$, the authors of [22] derived a converse bound under uncoded placement. They proved that the scheme has order optimality within a factor of 2. A transformation method was proposed in [23] to adapt the MN scheme to the multi-access scenario. This adaptation achieves a lower transmission load than [8] when $L\nmid K$. The work in [24] decomposed the multi-access caching problem into structured index coding problem. Based on solutions to the structured index coding problem, a new scheme has been proposed with a transmission load that does not exceed the transmission loads of the schemes in [22] and [23]. It is worth noting that subpacketization of the above scheme increases exponentially with KK.[25] proposed a consecutive cyclic placement strategy. By constructing a class of PDAs under this strategy, a novel multi-access coded caching scheme was introduced. This scheme has a subpacketization increasing linearly with the number of users and attains a minimal transmission load of $\frac{K-KML/N}{2\left\lfloor\frac{K}{K-KML/N+1}\right\rfloor+1}$. All the aforementioned schemes are designed for general system parameters. A summary is provided in Table I. For specific parameter configurations, several schemes with reduced subpacketization have been introduced in [26, 27, 28, 29]. These are also included in Table I.
TABLE I: Known multi-access coded caching schemes with uncoded placement.Where $K$ is the number of users, the number of cache nodes is equal to $K$, $L$ is the access degree and the cache-node memory ratio is $\frac{t}{K}$. If the range of $t$ is not specified in the Limitation section, then $t\in\{0,\ldots,\lfloor\frac{K}{L}\rfloor\}$.
@@
$\frac{1}{\left(\lceil\frac{2tL}{K-tL+1}\rceil+1\right)}+\sum_{r=\frac{K-tL+3}{2}}^{K-tL}\frac{2}{1+\lceil\frac{tL}{r}\rceil}$ $K-tL>1\text{ is odd};\text{$t$, $K$ are coprime}$
@@
Furthermore, the work in [30] introduces a multi-access coded caching scheme that incorporates secure delivery mechanisms. When the additional memory allocated for security purposes is excluded, this scheme achieves the same transmission load as the RK1 scheme presented in [22], and the required subpacketization level does not exceed $K$. In [31], a converse bound based on the cut-set argument was established, and a new scheme of coded placement was proposed for the memory regime satisfying M<(NK+L)/KM<(N-K+L)/K. This scheme is proven to be optimal under the condition $N\leq K$. Several studies [32, 33, 34] have investigated multi-access coded caching systems with a slightly modified topology based on cross-resolvable designs. The authors of [35, 36] explored a two-dimensional multi-access scenario, where the mobile users move on a two-dimensional grid and can access the nearest cache-nodes inside the grid. The work in [37] presents a novel scheme for a general MACC setting, unifying existing MACC schemes with the customized solution for specific encoding cache scenarios. This scheme enhances user distribution density while maintaining good load performance.
As observed in Table I, while the schemes in [8, 22, 23, 24] achieve competitive transmission loads, their subpacketization grows exponentially with the number of users. In contrast, [25, 28, 29] and other schemes maintain linear or polynomial subpacketization, but at the cost of limited coded caching gains and higher transmission loads. This trade-off reveals a key limitation in multi-access coded caching. Although the PDA serves as a combinatorial structure that characterizes placement and delivery strategies, its definition alone does not directly yield constructions that simultaneously achieve low subpacketization and significant coded caching gain. Therefore, constructing a general framework for PDA that guarantee linear subpacketization while maintaining near-optimal transmission performance is an important and meaningful research direction.
@@ I-B Research motivation and contribution
In this paper, we focus on designing multi-access coded caching schemes with linear subpacketization, aiming to minimum the transmission load under uncoded placement and one-shot delivery strategies. Through the analysis of Table I, we can find that existing multi-access coded caching schemes with linear or polynomial subpacketization still face challenges in balancing subpacketization and rate. Motivated by the structure of non-half-sum disjoint packing (NHSDP) in shared-link caching systems, we extend this interesting structure to the multi-access network and propose a novel structure called cyclic multi-access non-half-sum disjoint packing (CMA-NHSDP). This structure offers a unified combinatorial framework for designing multi-access coded caching schemes with linear subpacketization. More specifically, the main contributions are summarized as follows:
@@
  • We propose a novel combinational structure termed CMA-NHSDP. It is worth noting that the NHSDP proposed in [19] for shared-link networks cannot be directly applied to MACC systems, as each user in the shared-link system is equipped with independent caches. The structure CMA-NHSDP is a generalized framework for extending NHSDP to MACC systems. Given an $(L,v,g,b)$ CMA-NHSDP for any odd positive integers $v$, any positive integers $L$, $g$ and $b$, we can obtain an $L$-continuous (K=v,F=v,Z=vbg,S=bv)(K=v,F=v,Z=v-bg,S=bv) PDA. When $K$ is even, we can add one virtual user to construct a scheme based on the $(L,K+1,g,b)$ CMA-NHSDP.
@@ A $F$-division $(K,L,M,N)$ multi-access coded caching scheme contains two independent phases:
The transmission load (or rate) is defined as the worst case transmission amount $R=\max_{{\bf d}\in[1:N]^{K}}S_{\mathbf{d}}/F$. Our objective is to design a scheme such that the transmission load is as small as possible given the parameters $K$, $L$, $M$ and $N$.
@@
  • An $F\times K$ node-placement array $\mathbf{C}$ consists of stars and nulls, where $F$ and $K$ represent the subpacketization and the number of cache-nodes, respectively. The entry located at the position $(j,k)$ in $\mathbf{C}$ is star if and only if the $k^{\text{th}}$ cache-node caches the $j^{\text{th}}$ packet of each $\mathbf{w}_{n}$ where $n\in[N]$.
  • An $F\times K$ user-retrieve array $\mathbf{U}$ consists of stars and nulls, where $F$ and $K$ represent the subpacketization and the number of users, respectively. The entry at the position $(j,k)$ in $\mathbf{U}$ is star if and only if the $k^{\text{th}}$ user can retrieve the $j^{\text{th}}$ packet of each $\mathbf{w}_{n}$ where $n\in[N]$.
@@ II-B Placement delivery array
The authors in [23] also pointed out that if the user-delivery array $\mathbf{Q}$ satisfies the conditions C22 and C33, each user can decode its required file. That is the following statement. Specifically, if $\mathbf{Q}(j,k)=s$ is an integer, it implies that the $j^{\text{th}}$ packet of all files is not accessible to user $U_{k}$, and the server broadcasts a multicast message (i.e., the XOR of all the requested packets indicated by $s$) to all the users at time slot $s$. Condition C33 of Definition 2 ensures that each user can get its desired packet, since all the other packets in the multicast message are accessible to it. The occurrence number of integer $s$ in $\mathbf{Q}$, denoted by $g_{s}$, is called the coded caching gain of the PDA in time slot $s$, since the message broadcast at time slot $s$ can serve $g_{s}$ users simultaneously. Condition C2 of Definition 2 implies that the number of messages broadcasted by the server is exactly $S$, so the transmission load is $R=S/F$.
@@
    We can check that all the integers in $\mathbf{Q}$ satisfy Conditions C2 and C3 in Definition 2. According to Definition 1, we can obtain the following delivery strategy by $\mathbf{Q}$.
@@ III Cyclic Multi-access Non-half-sum disjoint packing
We can find that the stars occur exactly $Z=v-bg=27-3\times 4=15$ times in each column of $\mathbf{P}$, and appear in each row as $L=3$ cyclic consecutive. So the conditions C22 and C44 hold. In addition, there are exactly $S=81$ vectors in $\mathbf{P}$, and each vector occurs at most once in each row and each column. i.e., the conditions C22 and C33-a of Definition 2 hold. Finally, let us consider condition C33-b. Let us first consider the entries $\mathbf{P}(5,0)=\mathbf{P}0,5)=(5,3)$. We have $\mathbf{P}(5,5)=\mathbf{P}(0,0)=*$ which satisfies condition C33-b. Similarly, we can check that all the vectors in $\mathbf{P}$ satisfy condition C33-b) of Definition 2. Therefore, $\mathbf{P}$ is an $L$-continuous $(27,27,15,81)$ PDA that can realize a MACC scheme with memory ratio $\frac{M}{N}=\frac{5}{27}$, subpacketization $F=K=27$, and load $R=3$.
@@ IV Constructions of PDAs via CMA-NHSDP
By the above equation, $\mathfrak{D}$ has $m_{1}m_{2}m_{3}=3$ blocks and the intersection of any blocks is empty. The minimum and maximum values of $\mathfrak{D}$ are 40-40 and $40$, respectively. To ensure that these two sets $[40:0]$ and $[1:40]$ do not have overlap in $\mathbb{Z}_{v}$, the value of $v$ must satisfy $v\geq 2\times 40+1=81$, since 0 is also in $\mathbb{Z}_{v}$. Next, let us verify that $(\mathbb{Z}_{81},\mathfrak{D})$ is a $(3,81,8,3)$ CMA-NHSDP. Now let we consider the half-sum of $38$ and $16$, i.e., $\frac{38+16}{2}=27$, which is not in any block of $\mathfrak{D}$. Similarly, we can check the half-sum of any two different integers in each block of $\mathfrak{D}$. Finally, we can check that the sorted set $\mathscr{D}=\{\pm 14,\pm 15,\pm 16,\pm 20,\pm 21,\pm 22,\pm 32,\pm 33,\pm 34,\pm 38,\pm 39,\pm 40\}$ satisfies that its complement in $\mathbb{Z}_{7}$ can be partitioned into groups of $3$ cyclic consecutive integers, such as $\{17,18,19\}$.
@@
Next, given the parameters $v$ and $n$ (recall that $K=v$ and $g=2^{n}$), we consider how to select $m_{1},\ldots,m_{n}$ such that the memory ratio is minimized. That is, we aim to find the MACC scheme with the minimum required memory ratio while fixing the coded caching gain. Considering the constraint $v\geq 2\phi(m_{1},m_{2},\ldots,m_{n})+1$ into the optimization problem for the selection of $m_{1},m_{2},\ldots,m_{n}$, we have
@@
Since this equation (15) involves modulo arithmetic, it can not be solved directly. However, through abundant computer calculations, we have found that a good solution can be obtained when $m_{i}$ satisfies the following relationship.
@@
We now compare the performance of Theorem 4 and Construction 2 through a concrete example. Let $n=4$, $L=3$, and let $m_{2}$ vary in 1,2,3,4{1,2,3,4}. The corresponding parameters obtained from the two approaches are summarized in Fig. 2. It can be observed that Theorem 4 yields a smaller value of $v$ while maintaining the same objective function value $f$.
Figure 2: Tradeoff between $v$ and $f$